OpenStudy (anonymous):

Gretchen invested $1000 in a certificate of deposit(CD) that paid intrest at a rate of 9% per year. How much was the CD worth at the end of 3 years?

OpenStudy (tkhunny):

What have you tried? What are the compounding rules?

OpenStudy (anonymous):

well my teacher said i was suppost to muliply 1000 by the 9% after changing it to a decimal. then times that by 3 but the answer i got isnt in the answer choice.

OpenStudy (tkhunny):

First, let's fix the language. "times that by" doesn't mean anything. I sincerely hope your teacher didn't use that langauge. Second, what should have been communicated was to do the 9% thing three times. 1000*(1.09)*(1.09)*(1.09) This is annual interest and annual compounding.

OpenStudy (anonymous):

first of all my teacher didnt use that language and second its suppost to be like this ;\[1000 \times .09= 90.00 \] \[90.00 \times 3= 270\] but its not in the answer choices only 27 is but its already a answer to something else.

OpenStudy (tkhunny):

Simple Interest would be that, $270.00 in interest, and thus $1,270.00 total value at the end. Compound Interest 1000*1.09^3 = $1,295.03 It's probably supposed to be one of those. It's just not clear what is wanted. We need a better problem statement, but if that's all you have, we'll ahve to turn it back over to your teacher to explain.

OpenStudy (anonymous):

OK thank you very much its 1270. thank you!!!!

OpenStudy (tkhunny):

Well, then Simple Interest it is. The problem statement should tell us that. Good work!

OpenStudy (anonymous):

Yes its very hard for a 8th grader! (my teacher just did an example then gave us this work sheet.)